INAF LBC+LUCIFER queue observing, 2011 June 24 -25

Observer: R. Speziali, K. Boutsia
Telescope Operator: G. Bechetti
Support Astronomer: D. Thompson
Instrument: LUCIFER

Summary

Clear night with seeing ranging from 0.6'' to 0.9''.
We lost a total of 1h45min due to telescope and instrument problems: 30min because of lucifer (the mask mechanism got stack) and 1h15min because of the telescope (30min for pointing and collimation at the beginning of the night which is excessively long and 45min because rotator hit limits and we had to repeat a MOS acq).

Data Summary

Project Objects Filter ExpTime Notes
Turatto SN2011dh 200_H+K - OrderSep 2000sec arc+flat+tell BUT during the first offset the guide star was not reachable and the telescope was not guiding- for 500sec the object is not in the slit
Turatto SN2011dh H (img) 12sec  
Turatto SN2011dh J (img) 12sec  
Turatto SN2011dh Ks (img) 24sec  
PHOT_STD GD153 200_H+K - OrderSep 300sec requested by Turatto for calibration
Magrini 7c17_cluster 210_zJHK - H 3hr arc+flat + Tell (?) - the telluric was not visible when we acquired the field - after manually placing it in a slit we obtained 2 spectra

Details

02:11 - Initialize

02:37 - Twilight flats Ks,J,H
  • Ks : 0071-0090 (15000 - 9000cnts) / 0091-0100 (5000-6000cnts) / 0101-0110 (3000-4000cnts)
  • H : 0111-0120 (13000 - 10000cnts) / 0131-0140 (5000-3500cnts) / 0151-0160 (2000-1300cnts)
  • J : 0121-0130 (14000 - 10000cnts) / 0141-0150 (5000-3500cnts) / 0161-0170 (2000-1300cnts)

03:01 - Correct pointing

03:19 - STD_GD153_SpecAcquisitionTemplate.txt
  • error committing telescope - no guide star found
  • correct pointing again

03:24 - STD_GD153_SpecAcquisitionTemplate.txt again - failed
  • no guide star

03:27 - need to go back to a BS star because we are out of focus and this is why it does not find the guide star

03:37 - STD_GD153_SpecAcquisitionTemplate.txt 0171-0173
  • seeing on the guider 0.8''

03:46 - STD_GD153_SpecScienceTemplate 0174-0175
  • seeing on the guider 0.6''

03:53 - SN2011dh_SpecAcquisitionTemplate 0176-0178

04:07 - SN2011dh_SpecScienceTemplate 0179-0186
  • during the first offset the guide star was out of reach
  • for the first 1000sec of exposure we were not guiding
  • the SN can be seen in the first two images but NOT in the last 2 (0181-0182) - maybe they can be used as 'sky'
  • after it did the offset the guide star was found and in the rest of the images we can see the spectrum again

04:43 - STD_Tell_HD121409_SpecAcquisitionTemplate.txt 0187-0189
  • star was saturated - we decided to put a Br-gamma filter for the through-slit image
  • Tellurics best acquired with NB filters - note to be sent to PI

04:53 - STD_Tell_HD121409_SpecScienceTemplate.txt 0190-0191

04:55 - Calib Unit IN
  • ARC 0192-0199
  • FLAT 0200-0208

05:05 - Calib Unti OUT

05:06 - SN2011dh_ImageAcquisition 0209-0214

05:12 SN2011dh_ImageScience
  • J 0215-0218
  • Ks 0219-0226
  • H 0227-0230

05:21 - magrini_acq_sci.txt - MOS error!!! - preset failed
  • David has to recover the error
  • storage to turn-out - translator went to the mask- head rotation power on- grabbed the mask - unlock the mask
  • the error is known as 'disconnect error'

05:51 - Initialize

05:53 - magrini_acq_sci.txt 0231-0233
  • the starts for the MOS acq are too faint.
  • we need to repeat increasing exposure time

06:06 - magrini_acq_sci.txt 0234-0240
  • the new exp time is DIT=20 NDIT=6 and we also changed the filter to J so that the background is lower
  • after the MOS acquisition the mask still seems not well aligned
  • probably the PI should have put more alignment stars.
  • we try to apply rotation and offsets manually so that we can center the sources in the 2 square slits
  • final offsets applied PA=0.15deg, x,y=-0.24,0.24 and again PA=0.15deg
  • this whole procedure took us almost 50min

06:56 - magrini_sci_H_1.txt 0241-0247

07:34 - during the 8th offset the rotator hit a limit and the preset was canceled
  • 7 offsets executed
  • we need to repeat the MOS acq after the telescope gets unwrapped

07:50 - magrini_acq_sci.txt 0248-0252
  • we applied again the offsets manually in order to center the mask
  • almost 30 min lost because of rotator error!

08:16 - magrini_sci_H_2.txt 0253-0264
  • offset Y=+1.5 to center mask

09:07 - seeing on the guider is 1.0'' * DIMM 0.9''

09:24 - magrini_sci_H_3.txt 0265-0276
  • offset Y=+1.5

10:30 - magrini_sci_H_4.txt 0277-0282
  • lucifer gives temperature warnings...

11:04 - Calib unit ON -
  • ARC 0283-0290
  • FLAT 0291 - 0299

11:17 - Calib unit OUT

11:18 - 'magrini_tell_acq_PiHer.txt 0300-0304
  • we find no guiding star
  • correct pointing

11:29 - magrini_tell_acq_PiHer.txt - no guide star again
  • we take an image anyway to see if the star is in the field
  • we see no star in the field - only a ghost on the left side
  • we offset x,y=120,0 and we see the star in the lower left side
  • we take a through-slit image and we manually center the star in one slit (presumably the one that the PI intended) 0305-0310
  • we take some spectra but during the big offset the preset is canceled. In the first two frames we can see the telluric 0311-0312
  • the image 0313 is supposed to be the telluric after changing slit but we see nothing... preset canceled - image taken outside the script
  • the coordinates in the tell script do not correspond to PiHer as mentioned in the title
  • we had to do a huge manual offset in order to see the telluric - WHY??

11:52 - End of night - DoDarks script
  • dark frames 0314-0367

-- KonstantinaBoutsia - 25 Jun 2011
Topic revision: r5 - 26 Jun 2011, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback